OpenStudy (anonymous):

[1/(x+2)-(1/2)]/[x] Help! and show steps please

OpenStudy (anonymous):

so basically it looks like this

OpenStudy (anonymous):

do u need to solve for x?

OpenStudy (anonymous):

i need to simplify

OpenStudy (anonymous):

ok

OpenStudy (mertsj):

\[\frac{\frac{1}{x+2}-\frac{1}{2}}{x}\times \frac{2(x+2)}{2(x+2)}=\frac{2-(x+2)}{x(2)(x+2)}=\frac{-x}{2x(x+2)}=\frac{-1}{2(x+2)}\]

OpenStudy (anonymous):

why did you multiply it by 2(x+2)/2(x+2)?

OpenStudy (mertsj):

Because I wanted to cancel all the denominators.

OpenStudy (mertsj):

And that's what would cancel the denominators in the numerator so I had to multiply the denominator by the same thing.

OpenStudy (anonymous):

is that a general rule for cancelling denoimators? the rule being multiply numerator and denominator by the numerator's denominator?

OpenStudy (mertsj):

The general rule is multiply numerator and denominator by the common denominator of ALL denominators.

OpenStudy (mertsj):

Then all the denominators will cancel.

OpenStudy (anonymous):

ok i get it. thanks for your help :)

OpenStudy (mertsj):

yw